OpenStudy (anonymous):

Use a net to find the surface area of the right triangular prism shown below: Three rectangles next to each other with a width of 7 feet. The first and second rectangles length are unknown, and the last is 10 feet long. There are two right triangles above and below the last rectangle. One is labeled with a base of 6 feet and a height of 8 feet. 198 square feet 216 square feet 270 square feet 280 square feet

OpenStudy (anonymous):

@tabpernick @Snoopy_Lascano

OpenStudy (anonymous):

@mathmath333 @sammixboo @MTALHAHASSAN2 @kevindragonlord

OpenStudy (mathmath333):

is their a diagram given ?

OpenStudy (anonymous):

@mathmath333

OpenStudy (anonymous):

@AlexandervonHumboldt2

OpenStudy (mathmath333):



OpenStudy (mathmath333):

there are two similar right angled triangles and two similar rectangles and one distinct rectangle.

OpenStudy (anonymous):

ok but the awnswer is...? @mathmath333

OpenStudy (mathmath333):

\(\large \begin{align} \color{black}{ Total~Area=2\left(\dfrac{1}{2}\times 6\times 8\right)+7\times 26 \hspace{.33em}\\~\\ }\end{align}\) solve this